OpenStudy (anonymous):

a first number is 8 less than the second number. Twice the first number is 11 more than the second number. Find the numbers. im trying to figure out how to write this. i have x = y - 8, but im not sure how to write the second part, or if the first part is even correct.

OpenStudy (alexandervonhumboldt2):

a=b-8 2a=b+11

OpenStudy (alexandervonhumboldt2):

substitute: 2*(b-8)=b+11

OpenStudy (greatlife44):

Let's break this down first. A numbers 8 less than the second number. Let's say our second number we call it n. the first number is 8 less than that so it should be n-8. #1 2(n-8) #2 n+11

OpenStudy (greatlife44):

we would then set them equal to each other

OpenStudy (alexandervonhumboldt2):

2*(b-8)=b+11 2b-16=b+11 b=27

OpenStudy (alexandervonhumboldt2):

a=27-8=19

OpenStudy (greatlife44):

=)

OpenStudy (greatlife44):

the key here is that you must express both of these numbers using one variable

OpenStudy (greatlife44):

you had two variables in your proposed answer. that would be harder to do because you would need two equations to solve. but since we used one variable we had one equation and solved that to get the answer

OpenStudy (anonymous):

@greatlife44 , im supposed to be solving this as systems of equations, so in that case wouldnt i need two variables?